Marin Industries has the following patents on its December 31, 2019, balance sheet.

Patent Item

Initial Cost

Date Acquired

Useful Life at Date Acquired

Patent A

$41,208 3/1/16 17 years

Patent B

$15,240 7/1/17 10 years

Patent C

$15,360 9/1/18 4 years


The following events occurred during the year ended December 31, 2020.

1. Research and development costs of $232,000 were incurred during the year.
2. Patent D was purchased on July 1 for $30,438. This patent has a useful life of 91/2 years.
3. As a result of reduced demands for certain products protected by Patent B, a possible impairment of Patent B’s value may have occurred at December 31, 2020. The controller for Marin estimates the expected future cash flows from Patent B will be as follows.

Year

Expected Future Cash Flows

2021

$1,900

2022

1,900

2023

1,900


The proper discount rate to be used for these flows is 8%. (Assume that the cash flows occur at the end of the year.)

Compute the total carrying amount of Marin’ patents on its December 31, 2019, balance sheet.

Total carrying amount

$enter the Total carrying amount in dollars

Compute the total carrying amount of Marin' patents on its December 31, 2020, balance sheet.

Total carrying amount

$enter the Total carrying amount in dollars
Answer #1

1) Calculation of Carrying Amount on Dec 31, 2019 is shown as follows:-

Patent A

Useful Life in Months = 17 years*12 months = 204 months

Months used upto Dec 31, 2019 from Mar 1, 2016 = 2016+2017+2018+2019

= 10+12+12+12 = 46 months

Accumulated Amortization of Patent A upto Dec 31, 2019 = Cost of Patent A*(Month used/Useful Life in Months)

= $41,208*(46/204) = $9,292

Carrying Amount of Patent A on Dec 31, 2019 = Cost of Patent A - Accumulated Amortization

= $41,208 - $9,292 = $31,916

Patent B

Useful Life in Months = 10 years*12 months = 120 months

Months used upto Dec 31, 2019 from Jul 1, 2017 = 2017+2018+2019

= 6+12+12 = 30 months

Accumulated Amortization of Patent B upto Dec 31, 2019 = Cost of Patent B*(Month used/Useful Life in Months)

= $15,240*(30/120) = $3,810

Carrying Amount of Patent B on Dec 31, 2019 = Cost of Patent B - Accumulated Amortization

= $15,240 - $3,810 = $11,430

Patent C

Useful Life in Months = 4 years*12 months = 48 months

Months used upto Dec 31, 2019 from Sep 1, 2018 = 2018+2019

= 4+12 = 16 months

Accumulated Amortization of Patent C upto Dec 31, 2019 = Cost of Patent C*(Month used/Useful Life in Months)

= $15,360*(16/48) = $5,120

Carrying Amount of Patent C on Dec 31, 2019 = Cost of Patent C - Accumulated Amortization

= $15,360 - $5,120 = $10,240

Patent Carrying Amount on Dec 31, 2019
Patent A $31,916
Patent B $11,430
Patent C $10,240
Total $53,586

2) Calculation of Carrying Amount on Dec 31, 2020 is shown as follows:-

Patent A

Accumulated Amortization of Patent A upto Dec 31, 2020 = Cost of Patent A*(Month used/Useful Life in Months)

= $41,208*[(46+12)/204] = $11,716

Carrying Amount of Patent A on Dec 31, 2020 = Cost of Patent A - Accumulated Amortization

= $41,208 - $11,716 = $29,492

Patent B

Accumulated Amortization of Patent B upto Dec 31, 2020 = Cost of Patent B*(Month used/Useful Life in Months)

= $15,240*[30+12)/120] = $5,334

Carrying Amount of Patent B on Dec 31, 2020 = Cost of Patent B - Accumulated Amortization

= $15,240 - $5,334 = $9,906

Estimated Fair Value of Patent B = Expected Future Cash flows*PVAF(3 yrs, 8%)

= $1,900*2.577097 = $4,896

As estimated fair value of patent is less then its carrying amount on Dec 31, 2020, the patent B is impaired to its fair value (i.e. $4,896) Therefore the carrying amount of Patent B on Dec 31, 2020 is $4,896.

Patent C

Accumulated Amortization of Patent C upto Dec 31, 2020 = Cost of Patent C*(Month used/Useful Life in Months)

= $15,360*[(16+12)/48] = $8,960

Carrying Amount of Patent C on Dec 31, 2020 = Cost of Patent C - Accumulated Amortization

= $15,360 - $8,960 = $6,400

Patent D

Useful Life in Months = 9.5 years*12 months = 114 months

Months used upto Dec 31, 2020 from Jul 1, 2020 = 6 months

Accumulated Amortization of Patent D upto Dec 31, 2020 = Cost of Patent D*(Month used/Useful Life in Months)

= $30,438*[6/114] = $1,602

Carrying Amount of Patent D on Dec 31, 2020 = Cost of Patent D - Accumulated Amortization

= $30,438 - $1,602 = $28,836

Patent Carrying Amount on Dec 31, 2020
Patent A $29,492
Patent B $4,896
Patent C $6,400
Patent D $28,836
Total $69,624